# Fair Overlap Number of Balls (Fair-ONB): A Data-Morphology-based Undersampling Method for Bias Reduction

### v1

@@ 3.3 Structure and inner working of the Fair-ONB method
The Fair-ONB method is based on the characterisation of groups derived from the combination of class and protected feature values in the dataset. This is done via the data coverage using balls obtained from ONB.
Once the coverage is obtained, the Fair-ONB method eliminates instances of favoured groups to foster a fair classification. This is done in a guided way, selecting instances in problematic areas of the data space.
The Fair-ONB method selects the balls for preprocessing in two stages: first, the groups whose balls can be selected for elimination are chosen; then, using their characteristics, some balls of those groups are eliminated.
Group selection for preprocessing
When fairness metrics indicate bias towards a value of a protected feature, the chosen strategy is to perform undersampling of the groups with positive class (Ŷ=1) and said value (either 0 or 1) in the protected feature. Since there can be multiple protected features simultaneoulsy, there would be two ways of choosing the groups to preprocess.
@@
Depending on the dataset characteristics, the chosen undersampling strategy can greatly affect the results. For that reason, both group selection methods will be used.
Selection of balls from those groups for elimination
@@
  • Radius: if it is very small, the ball is in an area of group overlap, be it due to data distribution of for the existence of noise.
@@
  • Density: if it is very low, the ball might have been generated using an isolated point far from the decision boundary or just very far inside a class’ area, which might not contribute much to the model’s decisions. While some balls with very high density can also be problematic (due to, for example, having a very small radius), these particular cases are already covered by the other premises, so high density but useful balls will not be affected.
To perform the cleansing of the dataset, percentiles (values of an attribute for which exactly that percentage of cases have a lower value) of those three characteristics in the coverage are calculated, and thresholds for all three are chosen simultaneously. These thresholds are combined to decide different degrees of preprocessing (for example, one test can use percentile 5 for radius, 15 for number of covered instances and 10 for density). The balls from the selected preprocessing groups whose radius, number of covered instances or density are lower than each test’s chosen thresholds for each attribute are eliminated, along with the instances they cover, due to being considered problematic.
The utilisation of percentiles instead of predefined values aims to have a better adaptability to the characteristics of the different datasets.
@@ 4 Experimental Framework
This Section presents information regarding the experiments that test the goodness of the proposed method, Fair-ONB. The main aim is to check that, by using undersampling strategies guided by ONB, the fairness of models learned from the modified dataset can be improved while their performance is approximately maintained.
@@ 4.1 Datasets
This experimental framework includes the use of four datasets. The first three (COMPAS, Adult and German) are the most used datasets in the field of bias reduction; the fourth dataset, Ricci, was included to add more broadness to the study.
The exact versions of the aforementioned datasets have been modified to reduce the computational complexity: categorical variables were transformed into binary ones and, in the case of Adult, its size was cut in half (while maintaining its structure, via stratified sampling). The dataset characteristics (after those transformations) are as follows.
@@ 4.2 Experimental structure and parameters
    Testing the behaviour of the different threshold variables in different situations, to detect whether there are certain thresholds that generally give good results.
@@
    Comparing the bias reduction when using the Fair-ONB method to that of FAWOS, a recent neighbourhood-based sampling strategy.
@@
The parameters used for FAWOS are the same as in [18]. The weights for each label group are included in Table 2 (outliers always receive null weight), and oversampling factors were either 0.8, 1 or 1.2.
Table 2: Weights for each of FAWOS label groups.
Safe Borderline Rare
0 0.4 0.6
0 0.5 0.5
0 0.6 0.4
0.33 0.33 0.33
@@ 5.1 Study of the behaviour of threshold features in the Fair-ONB method
This Section presents the results of the experiments regarding the use of different threshold combinations. These experiments measure Disparate Impact according to the different protected variables of each dataset, as well as classifier performance according to AUC, for each different preprocessing parametrisation. Thus, the aims are to check whether fairness can be improved without much effect on AUC, and whether certain thresholding parameters can be widely considered useful.
@@
To facilitate a global comparison, Disparate Impact values for the different protected variables and AUC values are shown simultaneously; furthermore, to better reflect whether the method improves the results, the base results (with no preprocessing) are shown as a horizontal black line and the optimum Disparate Impact (1) as a green one.
Two sets of experiments are performed on each dataset, depending on whether union or intersection is used for undersampling group selection. In each case, for conciseness, radius will be used as the X axis feature, due to providing the most illustrative results.
Results on COMPAS
The undersampling results obtained using union undersampling are shown in Figure 3.
Figure 3: Plot of the behaviour of Disparate Impact according to the different protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using union undersampling on the COMPAS dataset.
@@
As can be observed, the results when different thresholds for the number of covered instances create result clusters regarding Disparate Impact, with noticeable differences. Meanwhile, an interesting behaviour is observed for intermediate values of radius thresholds, where Disparate Impact varies explosively. Generally, when using union undersampling, substantial Disparate Impact improvements can be attained on COMPAS (to the point of reaching the optimum on “sex”). AUC, while slightly reduced, is comparable to the base value.
The best combined disparate impact result (taking into account both protected features) was obtained using num_inst=1, radius=0.0015 and density=52.07, with a total distance to the Disparate Impact optima of 0.135 and a decrease of 0.009 in AUC.
As an example, Figure 4 shows two pie charts with the group distributions in the base dataset and when the aforementioned best performing undersampling is used. As can be observed, the best result is obtained without group balance.
Figure 4: Example of COMPAS group proportions in the base case (4a) and with the best-performing union undersampling (4b).
Regarding the intersection undersampling results, results are shown in Figure 5.
@@
Figure 5: Plot of the behaviour of Disparate Impact according to the different protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using intersection undersampling on the COMPAS dataset.
Just like in the union case, multiple clusters can be seen according to the number of covered instances threshold, and there is still a fast change in the intermediate values of the radius threshold. However, Disparate Impact is reduced less according to race.
The best global Disparate Impact result obtained when using intersection undersampling on COMPAS was obtained when using thresholds num_inst=2, radius=0.001250 and density=52.07, with a total distance to the optima of 0.133, and AUC only 0.007 below the base case.
As a summary of the results obtained with COMPAS, the best results of each type of undersampling (and the base case for comparison) are presented in Table 3.
Table 3: Comparison of Disparate Impact and AUC between the base COMPAS dataset and the best union and intersection undersamplings, along with the best undersamplings for the optimisation of each individual metric.
COMPAS Race DI Sex DI AUC
Baseline 0.837 1.086 0.654
Best Global Union DI 0.881 1.016 0.645
Best Race Union DI 0.965 0.680 0.643
Best Sex Union DI 0.853 1.000 0.649
Best Union AUC 0.874 0.805 0.662
Best Global Intersection DI 0.868 0.999 0.647
Best Race Intersection DI 0.885 0.966 0.643
Best Sex Intersection DI 0.866 0.999 0.647
Best Intersection AUC 0.831 1.089 0.656
Results on Adult
Following the same structure, the results using union undersampling on Adult are shown in Figure 6.
Figure 6: Plot of the behaviour of Disparate Impact according to the different protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using union undersampling on the Adult dataset.
On this dataset, union undersampling does not obtain noticeable improvements. Only slightly better Disparate Impact results were obtained for both protected variables. The best global Disparate Impact result when using union undersampling on Adult was obtained using num_inst=1, radius=0.0154 and density=0.88 as thresholds, with a total distance to the Disparate Impact optima of 0.890 and AUC 0.007 lower than the baseline.
The results obtained using intersection undersampling are the following (see Figure 7).
In this case, Disparate Impact provides considerable improvements for both protected variables, reaching the optimum in “race”. In the case of “sex”, much better results than the baseline are also obtained, while not reaching the recommended [0.8,1.25] Disparate Impact interval.
The best global Disparate Impact result obtained when using intersection undersampling on Adult was obtained when using thresholds num_inst=4, radius=0.0154 and density=6.73, with a total distance to the Disparate Impact optima of 0.384, and AUC 0.037 below the base case.
As before, Table 4 presents the summary of the best results on the Adult dataset.
@@
Figure 7: Plot of the behaviour of Disparate Impact according to the different protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using intersection undersampling on the Adult dataset.
Table 4: Comparison of Disparate Impact and AUC between the base Adult dataset and the best union and intersection undersamplings, along with the best undersamplings for the optimisation of each individual metric.
Adult Race DI Sex DI AUC
Baseline 0.675 0.379 0.746
Best Global Union DI 0.713 0.397 0.739
Best Race Union DI 0.723 0.363 0.730
Best Sex Union DI 0.713 0.397 0.739
Best Union AUC 0.675 0.379 0.746
Best Global Intersection DI 1.006 0.622 0.709
Best Race Intersection DI 0.995 0.619 0.709
Best Sex Intersection DI 0.983 0.622 0.710
Best Intersection AUC 0.675 0.379 0.746
Results on German
Likewise, the results using union undersampling on German are shown in Figure 8.
@@
Figure 8: Plot of the behaviour of Disparate Impact according to the different protected features (gender and age) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using union undersampling on the German dataset.
In this case, only slight improvements were obtained on “gender”, since it was already very close to the optimum Disparate Impact, while noticeable improvements were attained according to “age”. Regarding that first protected feature, the chosen number of covered instances threshold is very important, as increasing it leads to bad results.
The best global Disparate Impact result obtained when using union undersampling on German was obtained when using thresholds num_inst=1, radius=0.7109 and density=0.49, with a total distance to the Disparate Impact optima of 0.105, and AUC 0.013 over the baseline; thus, in this case it was possible to improve both fairness and performance simultaneously.
Regarding the intersection undersampling results, they are shown in Figure 9.
@@
Figure 9: Plot of the behaviour of Disparate Impact according to the different protected features (gender and age) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using intersection undersampling on the German dataset.
Employing intersection undestampling, a similar behaviour was observed (although less pronounced) when the number of covered instances threshold is modified, and the Disparate Impact ranges are lower. The best global Disparate Impact result obtained when using intersection undersampling on German was obtained when using thresholds num_inst=1, radius=0.5810 and density=0.44, with a total distance to the Disparate Impact optima of 0.096, and AUC only 0.015 below the base case.
The best results obtained on German are summarised in Table 5.
@@
Table 5: Comparison of Disparate Impact and AUC between the base German dataset and the best union and intersection undersamplings, along with the best undersamplings for the optimisation of each individual metric.
German Gender DI Age DI AUC
Baseline 1.028 1.295 0.633
Best Global Union DI 1.062 1.043 0.646
Best Gender Union DI 1.015 1.269 0.643
Best Age Union DI 1.062 1.043 0.646
Best Union AUC 1.139 1.270 0.599
Best Global Intersection DI 0.984 1.080 0.618
Best Gender Intersection DI 1.000 1.419 0.633
Best Age Intersection DI 0.984 1.080 0.618
Best Intersection AUC 1.229 1.229 0.649
Results on Ricci
Finally, regarding the Ricci dataset, there is only one protected feature (“race”), so it makes no sense to talk about union and intersection (since there is only one possible group to undersample). The undersampling results are shown in Figure 10.
Figure 10: Plot of the behaviour of Disparate Impact according to the race protected feature and of AUC when the radius threshold is modified and the other threshold variables are fixed, using undersampling on the Ricci dataset.
Slight improvements were obtained compared to the base case, as indicated in Table 6, obtaining the best global Disparate Impact result when using thresholds num_inst=1, radius=0 and density=5.49, with a distance to the Disparate Impact optimum of 0.487, and AUC 0.008 above the base case (so, in this case, both fairness and performance were simultaneously improved).
Table 6: Comparison of Disparate Impact and AUC between the base Ricci dataset and the best undersamplings according to each individual metric.
Ricci Race DI AUC
Baseline 0.501 0.908
Best Race DI 0.513 0.916
Best AUC 0.513 0.916
@@ 5.2 Sampling methods comparison
While the initial intention was to perform the comparison using all 4 datasets indicated in Section 4.1, the particularities of FAWOS would make using Adult unfeasible. In particular, the computational complexity of FAWOS, which is not prepared to work on datasets that big (even having already reduced its number of samples), is too high, given the fact that it uses 5NN for each sample multiple times per experiment. While the sample size could have been reduced again so that this dataset could have been added to this study, the necessary reduction would be so drastic that it would distort the dataset characteristics.
In this study, accuracy was used to measure classifier performance and, since FAWOS measures the Adapted Disparate Impact (see Equation 3 for reference), a simple transformation of Disparate Impact, that was the chosen fairness metric.
The best results obtained with each of the methods (and their optimal parameters) on the COMPAS, German and Ricci datasets are presented in Tables 7, 8 and 9, respectively. As a reminder, FAWOS parameters are the weights for safe (S), boderline (B) and rare (R) samples (as outliers receive weight 0), as well as the oversampling factor (OF); meanwhile, the parameters of the Fair-ONB method are the number of covered instances (N_i), radius (R) and density (D) thresholds.
@@
Fair-ONB Union Thresholds: N_i=3, R=0.00, D=2.47 0.886 0.972 0.142 0.771
Fair-ONB Intersec. Thresholds: N_i=1, R=0.00, D=34.37 0.873 0.999 0.128 0.747
@@
Fair-ONB Union Thresholds: N_i=1, R=0.57, D=0.36 0.942 0.786 0.262 0.672
Fair-ONB Intersec. Thresholds: N_i=1, R=0.80, D=0.22 0.925 0.789 0.286 0.682
@@
As can be observed, the Fair-ONB method obtains the best results according to Global Adapted Disparate Impact and, very often, also the best Adapted Disparate Impact on the individual protected features. Given the usual inverse relationship between predictive performance and bias mitigation, it makes sense that FAWOS obtains better accuracy results.

### v2

@@ 3.3 Structure and inner working of the Fair-ONB method
The Fair-ONB method is based on the characterisation of groups derived from the combination of class and protected feature values in the dataset. This is done via the data coverage using balls obtained from ONB (see Section 3.1).
Once the coverage is obtained, the Fair-ONB method undersamples instances of favoured groups to foster a fair classification. This is done in a guided way, selecting instances in problematic areas of the data space after harnessing the properties of coverage balls.
The Fair-ONB method selects the balls for undersampling in two stages: first, the groups whose balls can be selected for undersampling are chosen; then, using their characteristics, some balls of those groups (and the instances they contain) are eliminated.
Stage 1: group selection for preprocessing.
When Fairness metrics indicate bias towards a value of a protected feature in the dataset (which would likely translate into that same bias being learnt by the subsequent classifier), the chosen strategy is to perform undersampling of the groups with positive class (Ŷ=1) and said value (either 0 or 1) in the protected feature. Since there can be multiple protected features simultaneously, there would be two ways of choosing the groups to preprocess.
@@
Depending on the dataset characteristics, the chosen undersampling strategy can greatly affect the results. For that reason, both group selection methods will be studied.
Stage 2: selection of balls from those groups for elimination.
@@
  • Radius: if it is very small, the ball is in an area of group overlap, be it due to data distribution or to the existence of noise.
@@
  • Density: if it is very low, the ball might have been generated using an isolated point far from the decision boundary or just very far inside a class’ area, which might not contribute much to the model’s decisions. While some balls with very high density can also be problematic (due to, for example, having a very small radius), these particular cases are already covered by the other premises, so highly dense but useful balls will not be affected.
To perform the undersampling, percentiles (values of an attribute for which exactly that percentage of cases have a lower value) of those three characteristics in the coverage are calculated, and will serve as thresholds for ball selection. Using percentiles instead of predefined values aims to have a better adaptability to each dataset.
Once obtained, the thresholds of all three ball characteristics are combined to decide different degrees of preprocessing (for example, one test can simultaneously use percentile 5 for radius, 15 for number of covered instances and 10 for density). The balls from the selected preprocessing groups whose radii, number of covered instances or densities are lower than each test’s chosen ball characteristic thresholds are eliminated, along with the instances they cover, due to being considered problematic.
The overall modus operandi of the Fair-ONB method is shown in Figure 3.
Figure 3: The overall strategy of the Fair-ONB method starts with the generation of the groups defined by the pairs of class and protected feature values. The group coverage is then performed using ONB (in the plot, color is used to discern protected feature values, and instance symbols and circumference design to indicate class values). According to a chosen Fairness metric, groups are selected for undersampling. Finally, undersampling is performed, eliminating balls of that group employing thresholds (in the example, radius, density and number of covered instances were employed, respectively, from top to bottom).
@@ 4 Experimental Framework
This Section presents the experimental framework that tests the goodness of the proposed Fair-ONB method, with the aim of improving model Fairness while maintaining performance. Section 4.1 presents the datasets that will be used in the experiments. Section 4.2 explains the structure of the experiments and their parameters.
@@ 4.1 Datasets
This experimental framework includes the use of four datasets. The first three (COMPAS, Adult and German) are the most used in the field of bias reduction [5]; the fourth dataset, Ricci, was included to add more broadness to the study.
The exact versions of the aforementioned datasets have been modified to reduce the computational complexity: categorical variables were transformed into binary ones via one-hot encoding and, in the case of Adult, its size was cut in half (while maintaining its structure, via stratified sampling). The dataset characteristics (after those transformations) are as follows.
@@ 4.2 Experimental structure and parameters
    Testing the behaviour of the different threshold variables (according to their percentiles) in different situations, to detect whether there are certain thresholds that generally give good results.
@@
    Comparing the bias reduction when using the Fair-ONB method to that of FAWOS [22], a recent neighbourhood-based sampling strategy with whose strategy Fair-ONB shares some similarities.
@@
The parameters used for FAWOS are the same as in the original study [22]. The weights for each label group were 0 for safe, 0.6 for borderline and 0.4 for rare points (outliers always get null weight), and oversampling factors were either 0.8, 1 or 1.2.
@@ 5.1 Study of the behaviour of threshold features in the Fair-ONB method
This Section presents the results of the experiments on threshold combinations. Disparate Impact is measured according to the different protected variables of each of the four datasets, as well as classifier performance according to AUC, for each preprocessing parametrisation. Thus, the aims are to check whether Fairness can be improved without much effect on AUC, and whether certain thresholding parameters can be widely considered useful. Two sets of experiments are performed on each dataset, depending on whether union or intersection is used for undersampling group selection.
@@
To better exemplify the different observed behaviours of union and intersection undersampling, the results on COMPAS and Adult will be plotted. To facilitate a global comparison, in this type of figure, Disparate Impact values for both protected variables in the dataset and AUC values are shown simultaneously; furthermore, to better reflect whether the method improves the results, the base results (with no preprocessing) are shown as a horizontal black line and the optimum Disparate Impact (1) as a green one. In each case, radius will be used as the X axis feature, due to providing the most illustrative results.
COMPAS dataset.
@@
Figures 4 and 5 respectively depict the union and intersection undersampling results on this dataset. As can be observed, the results when using different thresholds for the number of covered instances create result clusters regarding Disparate Impact, with noticeable differences. Meanwhile, an interesting behaviour is observed for intermediate values of radius thresholds, where Disparate Impact varies greatly in this dataset. Substantial Disparate Impact improvements can be attained on COMPAS (to the point of reaching the optimum on “sex” with both types of undersampling, although with different parameters). AUC, while slightly reduced, is comparable to the base value.
@@
Figure 4: Behaviour of Disparate Impact according to the protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using union undersampling on COMPAS.
@@
Figure 5: Behaviour of Disparate Impact according to the protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using intersection undersampling on COMPAS.
Adult dataset.
Figures 6 and 7 show the union and intersection undersampling results on this dataset. Here, union undersampling does not obtain noticeable improvements. Only slightly better Disparate Impact results were obtained for both protected variables. However, with intersection undersampling, Disparate Impact provides considerable improvements for both protected variables, reaching the optimum in “race”. In the case of “sex”, much better results than the baseline are also obtained, while not reaching the often recommended [0.8,1.25] Disparate Impact interval.
@@
Figure 6: Behaviour of Disparate Impact according to the different protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using union undersampling on Adult.
@@
Figure 7: Behaviour of Disparate Impact according to the different protected features (race and sex) and of AUC when the radius threshold is modified and the other threshold variables are fixed, using intersection undersampling on Adult.
Other datasets.
For the German dataset, only slight improvements were obtained on “gender”, since it was already very close to the optimum Disparate Impact, while noticeable improvements were attained according to protected feature “age”. In this dataset, like in Adult, very similar global Disparate Impact results were obtained when using union and intersection undersamplings, and in both cases it was possible to improve both Fairness and performance simultaneously. As for the Ricci dataset, there is only one protected feature (“race”), so it makes no sense to separate union and intersection (since there is only one possible group to undersample). Slight improvements were obtained compared to the base case.
Lessons learned.
As a summary of the results obtained with all datasets, including the best results of each type of undersampling (and the base case for comparison) are presented in Table 2. For each dataset, the best results on each individual metric are presented in bold, while the results of the parametrisation that minimises DI globally (that it, the one that minimises the total distance to the optimum disparate impacts amongst the protected features) are presented in bold and italics.
@@
Table 2: Comparison of Disparate Impact and AUC between the base COMPAS, Adult, German and Ricci datasets and the best union and intersection undersamplings, along with the best undersamplings for the optimisation of each individual metric.
Dataset US Type Aim US Thresholds DI by protected feature Perf.
N_i R D Race DI Sex DI AUC
COMPAS Baseline - - - - 0.837 1.086 0.654
Union Global DI 1 0.0015 52.07 0.881 1.016 0.645
Race DI 2 0.0035 52.07 0.965 0.680 0.643
Sex DI 1 0.0015 2.41 0.853 1.000 0.649
AUC 2 0 34.95 0.874 0.805 0.662
Intersec Global DI 2 0.0013 52.07 0.868 0.999 0.647
Race DI 2 0.0035 52.07 0.885 0.966 0.643
Sex DI 2 0.0013 2.41 0.866 0.999 0.647
AUC 1 0 2.41 0.831 1.089 0.656
N_i R D Race DI Sex DI AUC
Adult Baseline - - - - 0.675 0.379 0.746
Union Global DI 1 0.0154 0.88 0.713 0.397 0.739
Race DI 1 0.0154 8.63 0.723 0.363 0.730
Sex DI 1 0.0154 0.88 0.713 0.397 0.739
AUC 1 0 0.88 0.675 0.379 0.746
Intersec Global DI 4 0.0154 6.73 1.006 0.622 0.709
Race DI 4 0 8.63 0.995 0.619 0.709
Sex DI 4 0 6.73 0.983 0.622 0.710
AUC 1 0 0.88 0.675 0.379 0.746
N_i R D Gender DI Age DI AUC
German Baseline - - - - 1.028 1.295 0.633
Union Global DI 1 0.7109 0.49 1.062 1.043 0.646
Gender DI 1 0.4383 0.22 1.015 1.269 0.643
Age DI 1 0.7109 0.49 1.062 1.043 0.646
AUC 1 0.7109 0.22 1.139 1.270 0.599
Intersec Global DI 1 0.5810 0.44 0.984 1.080 0.618
Gender DI 1 0.8257 0.36 1.000 1.419 0.633
Age DI 1 0.581 0.44 0.984 1.080 0.618
AUC 2 0.0363 0.44 1.229 1.229 0.649
N_i R D Race DI - AUC
Ricci Baseline - - - - 0.501 - 0.908
Union Race DI 1 0 5.49 0.513 - 0.916
AUC 1 0 5.49 0.513 - 0.916
All in all, according to the results shown in Figures 4 to 7, as well as the overall results in Table 2, the following conclusions can be reached.
  • The choice between union and intersection undersampling can greatly affect the minimisation of Disparate Impact, as shown by the results on Adult (where intersection was clearly the better choice). Given the fact that in the other cases both union and intersection fared similarly to one another, it would seem using intersection undersampling is the better choice overall.
  • Choosing the right number of covered instances threshold is always very important, as this threshold defines the different result clusters in the Disparate Impact plots. One of the extremes amongst the percentiles chosen (being either the lowest or highest value) gave the best results in all datasets, although which one is the right choice depends on the dataset.
  • The radius threshold shows the behaviour/shape of clusters defined by the number of covered instances threshold, whether it is mostly constant, increasing, decreasing, sigmoidal, etc. Depending on the location of the plotted cluster compared to the optimum Disparate Impact, different radius threshold values will be optimal.
  • The density threshold gives different heights in the results cluster, so using different threshold values helps fine-tune Disparate Impact depending on the clusters location compared to the optimum Disparate Impact.
@@ 5.2 Sampling methods comparison
While the initial intention was to perform the comparison using all 4 datasets indicated in Section 4.1, the particularities of FAWOS would make using Adult unfeasible. The computational complexity of FAWOS, which is not prepared to work on datasets that big (even having already reduced its number of samples), is too high, as it uses 5NN for each sample multiple times per experiment. While the sample size could have been reduced again to include the dataset in this study, the drastic reduction necessary would likely distort the dataset characteristics.
In this study, accuracy was used to measure classifier performance and, since FAWOS uses the Adapted Disparate Impact, which is a simple transformation of Disparate Impact (see Equation 3), that was the chosen Fairness metric.
The best results obtained with each method (and their optimal parameters) on the COMPAS, German and Ricci datasets are presented in Tables 3 to 5, respectively. As a reminder, FAWOS parameters are the weights for safe (S), boderline (B) and rare (R) samples (as outliers receive weight 0), as well as the oversampling factor (OF); meanwhile, the parameters of the Fair-ONB method are the number of covered instances (N_i), radius (R) and density (D) thresholds.
@@
Fair-ONB Intersec. Thresholds: N_i=3, R=0.00, D=2.47 0.886 0.972 0.142 0.771
Fair-ONB Union Thresholds: N_i=1, R=0.00, D=34.37 0.873 0.999 0.128 0.747
@@
Fair-ONB Intersec. Thresholds: N_i=1, R=0.57, D=0.36 0.796 0.942 0.262 0.672
Fair-ONB Union Thresholds: N_i=1, R=0.80, D=0.22 0.789 0.925 0.286 0.682
@@
As can be observed, the Fair-ONB method obtains the best results according to Global Adapted Disparate Impact (as indicated by the lower total distance to the optimum Adapted Disparate Impact amongst protected features) and, very often, also the best Adapted Disparate Impact on the individual protected features. Given the usual inverse relationship between predictive performance and bias mitigation, it makes sense that FAWOS obtains better accuracy results.
As a possible reason for this, Figure 8 shows two pie charts with the group distributions in the base COMPAS dataset and when the aforementioned best performing union undersampling is used. As can be observed, the best result is obtained without group balance, which is what most other methods (including FAWOS) would pursue, potentially giving Fair-ONB the edge. This also exemplifies the importance of correctly harnessing the morphology of protected groups towards modifying their proportions, as sampling in the most relevant parts of the dataset outweighed random sampling in group balancing (even if in FAWOS it was informed by neighbours according to 5NN).
Figure 8: Example of COMPAS group proportions in the base case (8a) and with the best-performing union undersampling using the Fair-ONB method (8b), showing that achieving group balance was not the optimal strategy.